OpenStudy (anonymous):

(3(x-2)^2)/(6(x-2)^4)

OpenStudy (anonymous):

\[\frac{1}{2(x-2)^2}\]

OpenStudy (mimi_x3):

= 3(x-2)(x+2) / 6(x-2)(x-2)(x+2)(x+2) = 1/2(x-2)(x+2) = 1/2(x-2)^2

OpenStudy (anonymous):

I agree!

OpenStudy (anonymous):

let the given expression be divided into two parts 3/6 * (x-2)^2/(x-2)^4 3/6 = 1/2 and (x-2)^2 = (x-2)(x-2) (x-2)^4 = (x-2)(x-2)(x-2)(x-2) \[3/6 * (x-2)^2/(x-2)^4\] \[(x-2)^2 = (x-2)(x-2) \] \[(x-2)^4 = (x-2)(x-2)(x-2)(x-2)\] hence \[[(x-2)(x-2)]/[(x-2)(x-2)(x-2)(x-2)]\] that is \[1/(x-2)(x-2)\] or \[1/(x-2)^2\] hence answer is \[1/[2 * (x-2)^2)]\]

OpenStudy (anonymous):

good

OpenStudy (anonymous):

got it :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

r u indian

OpenStudy (anonymous):

no

OpenStudy (anonymous):

then ?

OpenStudy (anonymous):

\[\color{indigo}{\text{hello}}\]